# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

### v2

@@ Acknowledgments
The author gratefully acknowledges the assistance of large language models in improving the grammar and style of this work.
Appendix
Appendix A Multilayer Perceptron and Backpropagation
@@
where $\mathbf{w}_{j}^{(1)}=[w_{j0}^{(1)},w_{j1}^{(1)},\ldots,w_{jN}^{(1)}]^{\mathrm{T}}$ represents the weight vector connecting the input layer to neuron $j$ in the first hidden layer $(l)$. Here, $w_{ji}^{(1)}$ denotes the weight connecting from the neuron $i$ in the input layer to the neuron $j$ in the first hidden layer. The output of neuron $j$ is then obtained by applying the activation function $g(\cdot)$:
@@
For subsequent hidden layers $(l=2,...,L)$, the activation $a_{j}^{(l)}$ for neuron $j$ in layer $l$ is computed as follows:
@@
where $\mathbf{z}^{(l)}=[z_{0}^{(l)},z_{1}^{(l)},\ldots,z_{M^{(l)}}^{(l)}]^{\mathrm{T}}\in\mathbb{R}^{M^{(l)}+1}$ is the output vector of layer $l$, $M^{(l)}$ is the number of neurons in this layer, and $z_{0}^{(l)}$ corresponds to the output of a bias neuron fixed at one. $w_{ji}^{(l)}$ denotes the weight connecting neuron $i$ in layer $l-1$ to neuron $j$ in layer $(l)$. The output for neuron $j$ is then given by
@@
This quantity, often referred to as the error for output neuron $k$, is the partial derivative of the loss $E(\mathbf{w})$ with respect to the neuron’s activation $a_{k}^{(L+1)}$. By chain rule, this derivative can be factored into two components: the partial derivative of the loss function with respect to the neuron’s output, $\frac{\partial E(\mathbf{w})}{\partial y_{k}}$, and and the derivative of the activation function evaluated at $a_{k}^{(L+1)}$, namely $f^{\prime}(a_{k}^{(L+1)})$.
To calculate the error $\delta_{j}^{(l)}$ for hidden layer $(l)$, we make use of the chain rule again. The partial derivative of the error function with respect to the activation $a_{j}^{(l)}$ is:
@@
The gradient of the error function with respect to the weights is computed using training data. In practice, to compute the gradient, most practitioners use a method called stochastic gradient descent (SGD) (LeCun et al., 2015). In SGD, the calculation of the gradient is performed on a subset of the training data, known as a mini-batch. This approach enables efficient computation of the gradient, making it feasible to train large neural networks using extensive datasets.

### v3

@@ Acknowledgments
The author gratefully acknowledges the assistance of Chat AI assistants (OpenAI ChatGPT and Google Gemini) and grammar checking service (Grammarly) in improving the grammar and style of this work.
Data Availability
The source code generated during this study is publicly available on GitHub at https://github.com/KazuhisaFujita/EDLA. The datasets analysed during this study are publicly available in the following repositories:
Funding Declaration
The author received no funding for this work.
Ethics declarations
Conflict of Interest
The author declares no competing interests.
Ethical Approval
This article does not contain any studies with human participants or animals performed by the author.
Appendix (for pedagogical completeness)
Appendix A Derivation of Backpropagation in a Multilayer Perceptron
@@
where $\mathbf{w}_{j}^{(1)}=[w_{j0}^{(1)},w_{j1}^{(1)},\ldots,w_{jN}^{(1)}]^{\mathrm{T}}$ represents the weight vector connecting the input layer to neuron $j$ in the first hidden layer $(l=1)$. Here, $w_{ji}^{(1)}$ denotes the weight connecting neuron $i$ in the input layer to neuron $j$ in the first hidden layer. The output of neuron $j$ is then obtained by applying the activation function $g(\cdot)$:
@@
For subsequent hidden layers (l=2,,L)(l=2,...,L), the activation $a_{j}^{(l)}$ for neuron $j$ in layer $l$ is computed as follows:
@@
where $\mathbf{z}^{(l-1)}=[z_{0}^{(l-1)},z_{1}^{(l-1)},\ldots,z_{M^{(l-1)}}^{(l-1)}]^{\mathrm{T}}\in\mathbb{R}^{M^{(l-1)}+1}$ is the output vector of layer l1l-1, $M^{(l-1)}$ is the number of neurons in this layer, and $z_{0}^{(l-1)}$ corresponds to the constant output of a bias neuron fixed to one. $w_{ji}^{(l)}$ denotes the weight connecting neuron $i$ in layer l1l-1 to neuron $j$ in layer $(l)$. The output for neuron $j$ is then given by
@@
This quantity, often referred to as the error for output neuron $k$, is the partial derivative of the loss $E(\mathbf{w})$ with respect to the neuron’s activation $a_{k}^{(L+1)}$. By the chain rule, this derivative can be factored into two components: the partial derivative of the loss function with respect to the neuron’s output, $\frac{\partial E(\mathbf{w})}{\partial y_{k}}$, and the derivative of the activation function evaluated at $a_{k}^{(L+1)}$, namely $f^{\prime}(a_{k}^{(L+1)})$.
To calculate the error $\delta_{j}^{(l)}$ for hidden layer $(l)$, we make use of the chain rule again. The partial derivative of the loss function with respect to the weight $w_{ji}^{(l)}$ is:
@@
The gradient of the loss function with respect to the weights is computed using training data. In practice, to compute the gradient, most practitioners use a method called stochastic gradient descent (SGD) [10]. In SGD, the calculation of the gradient is performed on a subset of the training data, known as a mini-batch. This approach enables efficient computation of the gradient, making it feasible to train large neural networks using extensive datasets.